### v3

# Treewidth 2 in the Planar Graph Product Structure Theorem

Abstract
We prove that every planar graph is contained in $H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$ for some graphs $H_{\_}1$ and $H_{\_}2$ both with treewidth 2. This resolves a question of Liu, Norin and Wood [arXiv:2410.20333]. We also show this result is best possible in the following sense: for any $c\in\mathbb{N}$, there is a planar graph $G$ such that for any tree $T$ and graph $H$ with $\operatorname{tw}(H)\leqslant 2$, $G$ is not contained in $H\boxtimes T\boxtimes K_{\_}c$.
keywords:
planar graph, product structure
1 Introduction
Graph product structure theory describes graphs in complicated graph classes as subgraphs of products of graphs in simpler graph classes, typically with bounded treewidth or bounded pathwidth. As defined in Section 2, the treewidth of a graph $G$, denoted by $\operatorname{tw}(G)$, is the standard measure of how similar $G$ is to a tree. As illustrated in Figure 1, the strong product $A\mathbin{\boxtimes}B$ of graphs $A$ and $B$ has vertex-set $V(A)\times V(B)$,
Figure 1: Strong product of paths.
where distinct vertices $(v,x),(w,y)$ are adjacent if:
  • $v=w$ and $xy\in E(B)$, or
  • $x=y$ and $vw\in E(A)$, or
  • $vw\in E(A)$ and $xy\in E(B)$.
The following Planar Graph Product Structure Theorem is the classical example of a graph product structure theorem. Here, a graph $H$ is contained in a graph $G$ if $H$ is isomorphic to a subgraph of $G$, written $H\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}G$.
Theorem 1.
For every planar graph $G$:
  1. (a)
    $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H\mathbin{\boxtimes}P$ for some graph $H$ with $\operatorname{tw}(H)\leqslant 6$ and path $P$ [21],
  2. (b)
    $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H\mathbin{\boxtimes}P\mathbin{\boxtimes}K_{\_}2$ for some graph $H$ with $\operatorname{tw}(H)\leqslant 4$ and path $P$ [21],
  3. (c)
    $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H\mathbin{\boxtimes}P\mathbin{\boxtimes}K_{\_}3$ for some graph $H$ with $\operatorname{tw}(H)\leqslant 3$ and path $P$ [8].
Dujmović et al. [8] first proved Theorem 1(a) with $\operatorname{tw}(H)\leqslant 8$. In follow-up work, Ueckerdt et al. [21] improved the bound to $\operatorname{tw}(H)\leqslant 6$. Part (b) is due to Dujmović, and is presented in [21]. Part (c) is in the original paper of Dujmović et al. [8]. Illingworth et al. [14] gave a new proof of part (c).
Theorem 1 provides a powerful tool for studying questions about planar graphs, by reducing to graphs of bounded treewidth. Indeed, this result has been the key for resolving several open problems regarding queue layouts [8], nonrepetitive colourings [7], centred colourings [5], adjacency labelling schemes [11, 2, 10, 6], twin-width [3, 17, 15], infinite graphs [13], and comparable box dimension [9]. In several of these applications, because the dependence on $\operatorname{tw}(H)$ is often exponential, the best bounds are obtained by applying the 3-term product in Theorem 1(c).
The $\operatorname{tw}(H)\leqslant 3$ bound in Theorem 1(c) is best possible in any result saying that every planar graph is contained $H\mathbin{\boxtimes}P\mathbin{\boxtimes}K_{\_}c$ where $P$ is a path (see [8]). Liu et al. [18] relaxed the assumption that $P$ is a path, and studied products of two graphs of bounded treewidth. They asked whether every planar graph is contained in $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$ for some graphs $H_{\_}1$ and $H_{\_}2$ with $\operatorname{tw}(H_{\_}1)\leqslant 2$ and $\operatorname{tw}(H_{\_}2)\leqslant 2$. We answer this question in the affirmative.
Theorem 2.
Every planar graph $G$ is contained in $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}2$ for some graphs $H_{\_}1$ and $H_{\_}2$ with $\operatorname{tw}(H_{\_}1)\leqslant 2$ and $\operatorname{tw}(H_{\_}2)\leqslant 2$.
We actually prove a strengthening of Theorem 2 that holds for a more general class of graphs $G$, and with a more precise statement about the structure of $H_{\_}1$ and $H_{\_}2$; see Theorem 7 below. We also show that Theorem 2 is best possible in the following sense.
Theorem 3.
For any integer $c\geqslant 1$ there is a planar graph $G$ such that for any tree $T$ and graph $H$ with $\operatorname{tw}(H)\leqslant 2$, $G$ is not contained in $H\mathbin{\boxtimes}T\mathbin{\boxtimes}K_{\_}c$.
We conclude this introduction by mentioning an open problem: Does Theorem 2 hold with $H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$ replaced by $H_{\_}1\boxtimes H_{\_}2$?
2 Treewidth
We consider finite simple undirected graphs $G$ with vertex-set $V(G)$ and edge-set $E(G)$. For a tree $T$ with $V(T)\neq\emptyset$, a $T$-decomposition of a graph $G$ is a collection $(B_{\_}x:x\in V(T))$ such that:
  • $B_{\_}x\subseteq V(G)$ for each $x\in V(T)$,
  • for every edge ${vw\in E(G)}$, there exists a node ${x\in V(T)}$ with ${v,w\in B_{\_}x}$, and
  • for every vertex ${v\in V(G)}$, the set $\{x\in V(T):v\in B_{\_}x\}$ induces a non-empty (connected) subtree of $T$.
The width of such a $T$-decomposition is ${\max\{|B_{\_}x|:x\in V(T)\}-1}$. A tree-decomposition is a $T$-decomposition for any tree $T$. A path-decomposition is a $P$-decomposition for any path $P$, denoted by the corresponding sequence of bags. The treewidth of a graph $G$, denoted $\operatorname{tw}(G)$, is the minimum width of a tree-decomposition of $G$. The pathwidth of a graph $G$, denoted $\operatorname{pw}(G)$, is the minimum width of a path-decomposition of $G$. By definition, $\operatorname{tw}(G)\leqslant\operatorname{pw}(G)$ for every graph $G$. Treewidth is the standard measure of how similar a graph is to a tree. Indeed, a connected graph has treewidth at most 1 if and only if it is a tree. It is an important parameter in structural graph theory, especially Robertson and Seymour’s graph minor theory, and also in algorithmic graph theory, since many NP-complete problems are solvable in linear time on graphs with bounded treewidth. See [12, 1, 19] for surveys on treewidth.
3 Proof of Theorem 2
Let $G^{+}$ be the graph obtained from a graph $G$ by adding one new vertex adjacent to every vertex in $G$. In any graph, a vertex $v$ is dominant if $v$ is adjacent to every other vertex. We use the following lemma by Liu et al. [18]. We include the proof for completeness.
Lemma 4 ([18]).
For any graph $G$ and any vertex-partition $\{V_{\_}1,V_{\_}2\}$ of $G$,
$G^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}G[V_{\_}1]^{+}\mathbin{\boxtimes}G[V_{\_}2]^{+}.$
Proof.
Let $Q:=G[V_{\_}1]^{+}\mathbin{\boxtimes}G[V_{\_}2]^{+}$, where $r_{\_}i$ is the dominant vertex in $G[V_{\_}i]^{+}$, for $i\in\{1,2\}$. Let $r$ be the dominant vertex in $G^{+}$. Map $r$ to $(r_{\_}1,r_{\_}2)$ in $Q$, which is adjacent to every other vertex in $Q$. Map each vertex $v\in V_{\_}1$ to $(v,r_{\_}2)$ in $Q$. Map each vertex $w\in V_{\_}2$ to $(r_{\_}1,w)$ in $Q$. For each edge $vv^{\prime}\in E(G[V_{\_}1])$, the images of $v$ and $v^{\prime}$ are adjacent in $Q$. For each edge $ww^{\prime}\in E(G[V_{\_}2])$, the images of $w$ and $w^{\prime}$ are adjacent in $Q$. For all vertices $v\in V_{\_}1$ and $w\in V_{\_}2$, the images of $v$ and $w$ are adjacent in $Q$. Hence, the above mapping shows that $G^{+}$ is isomorphic to a subgraph of $Q$. ∎
If every planar graph had a vertex-partition into two induced forests, then the Four-Colour Theorem would follow. Chartrand and Kronk [4] constructed planar graphs that have no vertex-partition into two induced forests. On the other hand, Thomassen [20, Theorem 4.1] showed the following analogous result where the forest requirement is relaxed [^1] Thomassen [20, Theorem 4.1] proved Lemma 5 for planar triangulations, which implies the result for general planar graphs, since any subgraph of a triangle-forest is a triangle-forest. Lemma 5 was rediscovered by Knauer et al. [16].. A triangle-forest is a graph in which every cycle is a triangle.
Lemma 5 ([20]).
Every planar graph has a vertex-partition into two induced triangle-forests.
We need the following elementary property of triangle-forests.
Lemma 6.
Every triangle-forest $G$ has a matching $M$ such that $G/M$ is a forest.
Proof.
We proceed by induction on $|V(G)|$. If $|V(G)|\leqslant 3$ then the claim holds trivially. Now assume that $|V(G)|\geqslant 4$. We may assume that $G$ is connected. Suppose that $\deg_{\_}G(v)=1$ for some $v\in V(G)$. By induction, $G-v$ has a matching $M$ such that $(G-v)/M$ is a forest. Since $\deg_{\_}G(v)=1$, $G/M$ is a forest. Now assume that $G$ has minimum degree at least 2. Let $B$ be a leaf block of $G$. Since $G$ has minimum degree at least 2, $B$ is 2-connected. If $B$ has two non-adjacent vertices $v$ and $w$, then a cycle through $v$ and $w$ has length at least four, contradicting that $G$ is a triangle-forest. So $B$ induces a triangle. Say $V(B)=\{u,v,w\}$ where $w$ is the cut-vertex separating $\{u,v\}$ from $G-V(B)$. By induction, $G-u-v$ has a matching $M$ such that $(G-u-v)/M$ is a forest. Hence $M^{\prime}:=M\cup\{uv\}$ is a matching in $G$ such that $G/M^{\prime}$ is a forest. ∎
Lemmas 6 and 5 imply:
Corollary 1.
Every planar graph $G$ has a matching $M$ such that $G/M$ has a vertex-partition into two induced forests.
For an integer $k\geqslant 0$, a graph $G$ is $k$-apex if $G-A$ is planar for some $A\subseteq V(G)$ with $|A|\leqslant k$. A graph $G$ is an apex-forest if $G-A$ is a forest for some $A\subseteq V(G)$ with $|A|\leqslant 1$. Every apex-forest has treewidth at most 2. Thus, the next result implies and strengthens Theorem 2.
Theorem 7.
Every $2$-apex graph $G$ is contained in $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}2$ for some apex-forests $H_{\_}1$ and $H_{\_}2$.
Proof.
We may assume that $G$ has an edge $ab$ such that $G^{\prime}:=G-a-b$ is planar. By Corollaries 1 and 4, $G^{\prime}$ has a matching $M^{\prime}$ such that $(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$ for some apex-forests $H_{\_}1$ and $H_{\_}2$. Thus, for the matching $M:=M^{\prime}\cup\{ab\}$ we have $G/M\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G^{\prime}/M^{\prime})^{+}\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2$. Hence $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}1\boxtimes H_{\_}2\boxtimes K_{\_}2$. Here we use the fact that if $M$ is a matching in a graph $G$, then $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}(G/M)\boxtimes K_{\_}2$. ∎
Note that an analogous proof shows that every $k$-apex graph $G$ is contained in $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}{\max\{k,2\}}$ for some apex-forests $H_{\_}1$ and $H_{\_}2$.
4 Proof of Theorem 3
To prove Theorem 3 it will be convenient to use the language of partitions. A partition $\mathcal{P}$ of a graph $G$ is a partition of $V(G)$, where each element of $\mathcal{P}$ is called a part. For a partition $\mathcal{P}$ of a graph $G$, let $G/\mathcal{P}$ be the graph obtained from $G$ by identifying the vertices in each non-empty part of $\mathcal{P}$ to a single vertex; that is, $V(G/\mathcal{P})$ is the set of non-empty parts in $\mathcal{P}$, where distinct parts $P_{\_}1,P_{\_}2\in\mathcal{P}$ are adjacent in $G/\mathcal{P}$ if and only if there exist $v_{\_}1\in P_{\_}1$ and $v_{\_}2\in P_{\_}2$ such that $v_{\_}1v_{\_}2\in E(G)$. A partition $\mathcal{P}$ of a graph $G$ is a tree-partition if $G/\mathcal{P}$ is contained in a tree, and $\mathcal{P}$ is a star-partition if $G/\mathcal{P}$ is contained in a star.
The next observation characterises when a graph is contained in $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$. We include the proof for completeness.
Observation 1 ([18]).
For any graphs $H_{\_}1,H_{\_}2$ and any $c\in\mathbb{N}$, a graph $G$ is contained $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$ if and only if $G$ has partitions $\mathcal{P}_{\_}1$ and $\mathcal{P}_{\_}2$ such that $G/\mathcal{P}_{\_}i\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}i$ for each $i\in\{1,2\}$, and $|A_{\_}1\cap A_{\_}2|\leqslant c$ for each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$.
Proof.
()($\Rightarrow$)( ⇒ ) Assume $G$ is contained $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$. That is, there is an isomorphism $\phi$ from $G$ to a subgraph of $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$. For each vertex $x\in V(H_{\_}1)$, let $A_{\_}x:=\{v\in V(G):\phi(v)\in\{x\}\times V(H_{\_}2)\times V(K_{\_}c)\}$. Similarly, for each vertex $y\in V(H_{\_}2)$, let $B_{\_}y:=\{v\in V(G):\phi(v)\in V(H_{\_}1)\times\{y\}\times V(K_{\_}c)\}$. Let $\mathcal{P}_{\_}1:=\{A_{\_}x:x\in V(H_{\_}1)\}$ and $\mathcal{P}_{\_}2:=\{B_{\_}y:y\in V(H_{\_}2)\}$, which are partitions of $G$. By construction, $G/\mathcal{P}_{\_}i\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}i$ for each $i\in\{1,2\}$. For $A_{\_}x\in\mathcal{P}_{\_}1$ and $B_{\_}y\in\mathcal{P}_{\_}2$, if $v\in A_{\_}x\cap B_{\_}y$ then $\phi(v)\in\{x\}\times\{y\}\times V(K_{\_}c)$. Thus $|A_{\_}x\cap B_{\_}y|\leqslant c$.
()($\Leftarrow$)( ⇐ ) Assume $G$ has partitions $\mathcal{P}_{\_}1$ and $\mathcal{P}_{\_}2$ such that $G/\mathcal{P}_{\_}i\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H_{\_}i$ for each $i\in\{1,2\}$, and $|A_{\_}1\cap A_{\_}2|\leqslant c$ for each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$. Let $\phi_{\_}i$ be an isomorphism from $G/\mathcal{P}_{\_}i$ to a subgraph of $H_{\_}i$. For each $A_{\_}1\in\mathcal{P}_{\_}1$ and $A_{\_}2\in\mathcal{P}_{\_}2$, enumerate the at most $c$ vertices in $A_{\_}1\cap A_{\_}2$, and map the $i$-th such vertex to $(\phi_{\_}1(A_{\_}1),\phi_{\_}2(A_{\_}2),i)$. This defines an isomorphism from $G$ to a subgraph of $H_{\_}1\mathbin{\boxtimes}H_{\_}2\mathbin{\boxtimes}K_{\_}c$, as desired. ∎
As illustrated in Figure 2, a graph $F$ is a fan if $F$ has a dominant vertex $v$, called the centre of $F$, such that $F-v$ is a path. A graph $F$ is a double-fan if $F$ has two dominant vertices $v$ and $w$, called the centres of $F$, such that $F-v-w$ is a path. Note that every double-fan is a planar triangulation.
(a) Refer to caption   (b) Refer to caption
Figure 2: (a) fan, (b) double-fan
Throughout the following proofs, we use the following convention: if $\mathcal{P}$ and $\mathcal{Q}$ are partitions of a graph $G$, and $v_{\_}i\in V(G)$, then let $P_{\_}i$ be the part of $\mathcal{P}$ with $v_{\_}i\in P_{\_}i$, and let $Q_{\_}i$ be the part of $\mathcal{Q}$ with $v_{\_}i\in Q_{\_}i$. Of course, it is possible that $P_{\_}i=P_{\_}j$ or $Q_{\_}i=Q_{\_}j$ for distinct vertices $v_{\_}i,v_{\_}j$.
Lemma 8.
For any $c\in\mathbb{N}$, let $F$ be a fan on at least $c^{2}+c+1$ vertices with centre $v_{\_}1$. Let $\mathcal{P},\mathcal{Q}$ be partitions of $F$ such that $\mathcal{P}$ is a tree-partition and $|P\cap Q|\leqslant c$ for all $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$. Then there exists $v_{\_}2\in V(F-v_{\_}1)$ such that $Q_{\_}1\neq Q_{\_}2$.
Proof.
Let $B$ be the path $F-v_{\_}1$. Since $v_{\_}1$ is dominant in $F$, and $\mathcal{P}$ is a tree-partition, $\mathcal{P}$ is a star-partition with centre $P_{\_}1$. Since $|P_{\_}1\cap Q_{\_}1|\leqslant c$ and $v_{\_}1\in P_{\_}1\cap Q_{\_}1$, we have $|V(B)\cap Q_{\_}1\cap P_{\_}1|\leqslant c-1$. Thus $B-(Q_{\_}1\cap P_{\_}1)$ has at most $c$ components. Since $|V(B)|>(c-1)+c^{2}$, there is a path component $B^{\prime}$ of $B-(Q_{\_}1\cap P_{\_}1)$ on at least $c+1$ vertices. If $B^{\prime}$ contains a vertex $v_{\_}2\in P_{\_}1$, then $v_{\_}2\notin Q_{\_}1$, as desired. So we may assume that $B^{\prime}$ does not intersect $P_{\_}1$. Thus $B^{\prime}$ is contained in a single part $P_{\_}2\in\mathcal{P}$. Since $|P_{\_}2\cap Q_{\_}1|\leqslant c<|V(B^{\prime})|$, there exists $v_{\_}2\in B^{\prime}$ such that $v_{\_}2\notin Q_{\_}1$, as desired. ∎
Lemma 9.
For any $c\in\mathbb{N}$, let $F$ be a double-fan on at least $18c^{2}+2c+$ 1 vertices. Let $\mathcal{P},\mathcal{Q}$ be partitions of $F$ such that $\mathcal{P}$ is a tree-partition and $|P\cap Q|\leqslant c$ for all $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$. Let $v_{\_}1,v_{\_}2$ be the centres of $F$. If $P_{\_}1\neq P_{\_}2$ and $Q_{\_}1\neq Q_{\_}2$, then there exist vertices $v_{\_}3,v_{\_}4$ such that $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ is a $4$-clique and $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ are pairwise distinct.
Proof.
Let $B$ be the path $F-v_{\_}1-v_{\_}2$. Since $v_{\_}1$ and $v_{\_}2$ are both adjacent to every vertex in $B$, and $v_{\_}1$ and $v_{\_}2$ are in distinct parts of $\mathcal{P}$, and since $\mathcal{P}$ is a tree-partition, $V(F)=P_{\_}1\cup P_{\_}2$. Since $|P\cap Q|\leqslant c$ for all $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$, this means that $|Q|\leqslant 2c$ for each $Q\in\mathcal{Q}$. In particular, since $v_{\_}1\in Q_{\_}1$ and $v_{\_}2\in Q_{\_}2$, this means that $|B\cap(Q_{\_}1\cup Q_{\_}2)|\leqslant 4c-2$. Thus $B-(Q_{\_}1\cup Q_{\_}2)$ has at most $14c-$ 1 components. Since $|V(B)|\geqslant 8c^{2}+2c-1>(4c-2)+(4c-1)2c$, some path component $B^{\prime}$ of $B-(Q_{\_}1\cup Q_{\_}2)$ has at least $12c+$ 1 vertices. Since $|Q|\leqslant 2c$ for each $Q\in\mathcal{Q}$, $B^{\prime}$ intersects at least two parts of $\mathcal{Q}$. In particular, there is an edge $v_{\_}3v_{\_}4$ of $B^{\prime}$ such that $Q_{\_}3\neq Q_{\_}4$. By the choice of $B^{\prime}$, we have $\{Q_{\_}3,Q_{\_}4\}\cap\{Q_{\_}1,Q_{\_}2\}=\emptyset$. Thus $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ is the desired $4$-clique. ∎
A distension of a graph $G$ is any graph $\widehat{G}$ obtained from $G$ by adding, for each edge $vw$ of $G$, a path $P_{\_}{vw}$ complete to $\{v,w\}$, where $P_{\_}{vw}\cap G=\emptyset$ and $P_{\_}{vw}\cap P_{\_}{ab}=\emptyset$ for distinct $vw,ab\in E(G)$. Here ‘complete to’ means that each vertex in $P_{\_}{vw}$ is adjacent to both $v$ and $w$. Note that $\widehat{G}[V(P_{\_}{vw})\cup\{v,w\}]$ is a double-fan, which we denote by $\widehat{G}_{\_}{vw}$. We say $\widehat{G}$ is the $t$-distension of $G$ if $|V(P_{\_}{vw})|=t$ for each edge $vw\in E(G)$. Observe that every distension of a planar graph is planar.
Lemma 10.
For any $c\in\mathbb{N}$, there exists a planar graph $G$ such that for any tree-partition $\mathcal{P}$ of $G$ and for any partition $\mathcal{Q}$ of $G$ with $|P\cap Q|\leqslant c$ for each $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$, there is a 4-clique $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ in $G$ such that $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ are pairwise distinct.
Proof.
Let $t:=8c^{2}+2c-1$. Let $F$ be the fan on $t+1$ vertices with centre $v_{\_}1$. Let $B$ be the path $F-v_{\_}1$. Let $J$ be the $t$-distension of $F$, and let $G$ be the $t$-distension of $J$. Since $F$ is planar, $J$ is planar, and $G$ is planar. Consider any tree-partition $\mathcal{P}$ of $G$ and any partition $\mathcal{Q}$ of $G$ with $|P\cap Q|\leqslant c$ for each $P\in\mathcal{P}$ and $Q\in\mathcal{Q}$.
Since $|V(F)|=t+1\geqslant c^{2}+c+1$, by Lemma 8 applied to $F$ and the induced partitions of $F$, there exists $v_{\_}2\in B$ such that $Q_{\_}1\neq Q_{\_}2$. Let $C$ be the $t$-vertex path $J_{\_}{v_{\_}1v_{\_}2}-v_{\_}1-v_{\_}2$.
Since $|V(J_{\_}{v_{\_}1v_{\_}2})|=t+2=8c^{2}+2c+1$, if $P_{\_}1\neq P_{\_}2$ then by Lemma 9, there exist vertices $v_{\_}3,v_{\_}4$ such that $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ is a $4$-clique $J_{\_}{v_{\_}1v_{\_}2}$ with $Q_{\_}1,Q_{\_}2,Q_{\_}3,Q_{\_}4$ pairwise distinct, as desired. So we may assume that $P_{\_}1=P_{\_}2$.
Consider any vertex $v^{\prime}\in V(C)$. Let $P^{\prime}\in\mathcal{P}$ and $Q^{\prime}\in\mathcal{Q}$ such that $v^{\prime}\in P^{\prime}\cap Q^{\prime}$. Note that $v^{\prime}v_{\_}1$ and $v^{\prime}v_{\_}2$ are edges of $J$, so the double-fans $G_{\_}{v^{\prime}v_{\_}1}$ and $G_{\_}{v^{\prime}v_{\_}2}$ exist. Since $Q_{\_}1\neq Q_{\_}2$, there exists $i\in\{1,2\}$ such that $Q^{\prime}\neq Q_{\_}i$. If $P^{\prime}\neq P_{\_}1$ (and thus $P^{\prime}\neq P_{\_}2$), since $|V(G_{\_}{v^{\prime}v_{\_}i})|=t+2=8c^{2}+2c+1$, by Lemma 9 there exists a 4-clique $\{v_{\_}i,v^{\prime},v_{\_}3,v_{\_}4\}$ in $G_{\_}{v^{\prime}v_{\_}i}$ with $Q_{\_}i,Q^{\prime},Q_{\_}3,Q_{\_}4$ pairwise distinct, as desired.
So we may assume that $V(C)\subseteq P_{\_}1$. Since $v_{\_}i\in P_{\_}1\cap Q_{\_}i$ for $i\in\{1,2\}$, and since $|P_{\_}1\cap Q|\leqslant c$ for all $Q\in\mathcal{Q}$, we have $|C\cap(Q_{\_}1\cup Q_{\_}2)|\leqslant 2c-2$. So $C-(Q_{\_}1\cup Q_{\_}2)$ has at most $12c-$ 1 components. Since $|V(C)|=t=8c^{2}+2c-1>(2c-1)c+(2c-2)$, there exists a path component $C^{\prime}$ of $C-(Q_{\_}1\cup Q_{\_}2)$ with at least $c+1$ vertices. Note that $C^{\prime}\subseteq P_{\_}1$ also. Since $|P_{\_}1\cap Q|\leqslant c$ for all $Q\in\mathcal{Q}$, there is an edge $v_{\_}3v_{\_}4$ in $C^{\prime}$ such that $Q_{\_}3\neq Q_{\_}4$. By the choice of $C^{\prime}$, $\{Q_{\_}3,Q_{\_}4\}\cap\{Q_{\_}1,Q_{\_}2\}=\emptyset$. So $\{v_{\_}1,v_{\_}2,v_{\_}3,v_{\_}4\}$ is the desired $4$-clique in $G$. ∎
The next result follows from Lemmas 10 and 1, which implies Theorem 3.
Theorem 11.
For any $c\in\mathbb{N}$ there exists a planar graph $G$ such that for any tree $T$ and graph $H$, if $G\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H\boxtimes T\boxtimes K_{\_}c$ then $K_{\_}4\mathrel{\begin{subarray}{c}\textstyle\subset\\[-2.58334pt] \textstyle\sim\\[-1.72218pt] \end{subarray}}H$ and $\operatorname{tw}(H)\geqslant 3$.
Theorem 11 strengthens a result of Dujmović et al. [8], who proved it when $T$ is a path.
The next lemma implies that the graph $G$ in Theorem 11 has bounded treewidth and bounded pathwidth. In particular, since $G$ is a distension of a distension of a fan $F$, and since $\operatorname{tw}(F)\leqslant\operatorname{pw}(F)\leqslant 2$, we have $\operatorname{tw}(G)\leqslant 3$ and $\operatorname{pw}(G)\leqslant\operatorname{pw}(\widehat{F})+2\leqslant\operatorname{pw}(F)+4\leqslant 6$. (With a more detailed analysis, one can prove Theorem 11 with $\operatorname{pw}(G)\leqslant 4$ for a slightly different graph $G$; we omit this result.)
Lemma 12.
For any graph $G$ and any distension $\widehat{G}$ of $G$,
$\operatorname{tw}(\widehat{G})\leqslant\max\{\operatorname{tw}(G),3\}\quad\text{and}\quad\operatorname{pw}(\widehat{G})\leqslant\operatorname{pw}(G)+2.$
Proof.
We first prove the treewidth bound. Consider a tree-decomposition $(B_{\_}x:x\in V(T))$ of $G$ with width $\operatorname{tw}(G)$. Apply the following operation for each edge $vw\in E(G)$. Let $x_{\_}0\in V(T)$ such that $v,w\in B_{\_}{x_{\_}0}$. Say $P_{\_}{vw}=(u_{\_}1,\dots,u_{\_}t)$ is the path complete to $\{v,w\}$ in $\widehat{G}$. Add new vertices $x_{\_}1,\dots,x_{\_}{t-1}$ and edges $x_{\_}0x_{\_}1,x_{\_}1x_{\_}2,\dots,x_{\_}{t-2}x_{\_}{t-1}$ to $T$. Let $B_{\_}{x_{\_}i}:=\{v,w,u_{\_}i,u_{\_}{i+1}\}$ for each $i\in\{1,\dots,t-1\}$. We obtain a tree-decomposition of $\widehat{G}$, in which every new bag has size 4. Thus $\operatorname{tw}(\widehat{G})\leqslant\max\{\operatorname{tw}(G),3\}$.
We now prove the pathwidth bound. Let $(B_{\_}1,\dots,B_{\_}m)$ be a path-decomposition of $G$ with width $\operatorname{pw}(G)$. By duplicating bags, we may assume there is an injection $f:E(G)\to\{1,\dots,m\}$ such that $v,w\in B_{\_}{f(vw)}$ for each edge $vw\in E(G)$. For each edge $vw\in E(G)$, if $i:=f(vw)$ and $P_{\_}{vw}=(u_{\_}1,\dots,u_{\_}t)$ as above, then insert the sequence of bags $B_{\_}i\cup\{u_{\_}1,u_{\_}2\},B_{\_}i\cup\{u_{\_}2,u_{\_}3\},\dots,B_{\_}i\cup\{u_{\_}{t-1},u_{\_}t\}$ between $B_{\_}i$ and $B_{\_}{i+1}$. Since $f$ is an injection, we obtain a path-decomposition of $\widehat{G}$, in which each bag has size at most two more than the original bag. Thus $\operatorname{pw}(\widehat{G})\leqslant\operatorname{pw}(G)+2$. ∎
Acknowledgements
Distel is supported by an Australian Government Research Training Program Scholarship. Hendrey and Wood are supported by the Australian Research Council. Karol and Yip are supported by Monash Graduate Scholarships.